from Jacqui and me to this record of our experiences as new beekeepers in The Marches.

Search This Blog

Friday, 22 January 2010

varroa monitoring

Today the temperature has reached 9c. Mite drop count 142 making the total for 4 days of 1035. Today there were some wax flakes on the floor indicating some feeding activity in the hive. Housekeeping continues with 9 dead bees by the entrance. Steady rain/drizzle has kept the bees inside.

No comments:

Post a Comment